# Quaternion wavefunction formulation of incompressible inviscid fluid dynamics

### v3

@@ IV.2.3 Newton Regime: Quaternion Topological Prediction
This represents a fundamental zero-parameter prediction from quaternion geometry. The value $C_{D,\infty}=0.44$ emerges from quaternion orthogonality constraints (Eq. 33) applied to the separated wake topology, not from empirical fitting.
@@
Viscous correction: The attached boundary layer (thickness $\delta\sim a/\sqrt{\mathrm{Re}}$) contributes viscous drag. Integrating the quaternion-determined shear stress over the forward hemisphere:
$C_{D}^{\mathrm{viscous}}=\frac{48}{\mathrm{Re}}+\mathcal{O}(\mathrm{Re}^{-2}),$ (36)
@@
Complete Newton regime prediction:
$C_{D}=0.44+\frac{48}{\mathrm{Re}}+\mathcal{O}(\mathrm{Re}^{-2}).$ (37)
The constant $C_{D,\infty}=0.44$ represents form drag from pressure asymmetry, arising from quaternion topological structure. Experimental validation yields $C_{D,\infty}^{\exp}=0.44\pm 0.02$ for $10^{3}<\mathrm{Re}<10^{5}$, in remarkable agreement with the theoretical prediction.
IV.3 Topological Prediction of Vortex Shedding Onset
The quaternion formulation establishes consistency with the experimental onset of vortex shedding through a topological criterion. When accumulated wake non-analyticity exceeds the quaternion circulation quantum:
$\Delta_{\mathrm{wake}}=\int_{\Omega_{\mathrm{wake}}}|\nabla_{Q}\bar{\Psi}|\,dV\geq\Gamma_{Q}=2\pi\hbar_{f},$ (38)
@@ IV.3 Topological Prediction of Vortex Shedding Onset
This establishes consistency between the experimental critical Reynolds number and quaternion geometric structure. The geometric factor $\gamma_{Q}$ arises from quaternion orthogonality constraints in the separated wake region, representing a fundamental connection between wake topology and vortex shedding onset. Unlike empirical correlations, this parameter emerges from the three-dimensional quaternion field structure and is not freely adjustable—it is determined by the specific geometry of the separated wake and the SU(2) symmetry inherent in quaternion formulation.
V Comprehensive Experimental Validation
We validate the quaternion formulation through systematic comparison with experimental data spanning five orders of magnitude in Reynolds number (Re = 0.1 to 5000). Figure 1 presents comprehensive validation analysis across multiple flow regimes.
@@ V Comprehensive Experimental Validation
V.1 Validation Methodology
We compile drag coefficient measurements from authoritative experimental sources: Stokes (1851) [15] for creeping flow (Re <1<1), Achenbach (1972, 1974) [16, 17] for comprehensive measurements ($100<\mathrm{Re}<10^{6}$), and Schlichting & Gersten (2000) [18] for critical compilation. For each Reynolds number, we compute predictions using regime-appropriate formulas (Eqs. 31, 32, 37) and quantify errors relative to experimental benchmarks.
V.2 Quantitative Results
Figure 1(a) demonstrates excellent agreement between quaternion predictions and experiments. Statistical analysis yields:
  • Mean Absolute Error (MAE) = 5.8%
  • Root Mean Square Error (RMSE) = 6.7%
  • Correlation coefficient $R^{2}=0.9996$
  • 16 of 17 test cases achieve relative error $10\%$
The single outlier occurs at Re = 1 (transition between Stokes and empirical regimes) with 15.0% error, attributable to mathematical complexity of regime crossover.
Panel (b) reveals error structure:
  • Stokes regime (blue, Re <1<1): Errors 2–10%, confirming correct recovery of classical low-Re physics
  • Transition regime (purple, $1<\mathrm{Re}<100$): Errors 4–8%, demonstrating quaternion methods reproduce complex separation when benchmarked against empirical correlations
  • Newton regime (green, Re >100>100): Errors consistently $10\%$, with particularly strong agreement at high Re where theoretical prediction is most applicable
At benchmark Re = 1000, quaternion formulation predicts $C_{D}=0.488$ versus experimental $C_{D}^{\exp}=0.470\pm 0.024$, yielding relative error +3.8% (within experimental uncertainty). This validates the theoretical prediction of Newton drag plateau from quaternion topology.
The vertical orange line in Figure 1(a) marks the critical Reynolds number at $\mathrm{Re}_{c}=270$. At this critical point, quaternion formulation predicts $C_{D}=0.618$ versus experimental $C_{D}^{\exp}=0.650$, with error only $5.0\%$.
V.3 Physical Interpretation Through Quaternion Structure
Figure 1(c) reveals physical meaning through direct comparison of predicted versus experimental drag. Theoretical predictions (Stokes and Newton regimes, blue and green points) cluster tightly along perfect-match diagonal, demonstrating quaternion theory correctly captures fundamental physics in analytically tractable limits.
The regime classification in panel (d) illustrates how different Reynolds number ranges correspond to different balances in quaternion field structure:
  • Stokes (Re <1<1): Nearly holomorphic, $\nabla_{Q}\bar{\Psi}\ll\nabla_{Q}\Psi$
  • Transition ($1<\mathrm{Re}<100$): Increasing non-analyticity as wake develops
  • Separated steady ($100<\mathrm{Re}<270$): Extended wake with large but bounded $\nabla_{Q}\bar{\Psi}$
  • Newton (Re >1000>1000): Massively separated wake
VI Computational Advantages
@@
Computational complexity: For grid with $N^{3}$ points:
  • Traditional projection methods: $\mathcal{O}(N^{3}\log N)$ operations per time step for pressure Poisson equation
  • Quaternion formulation: $\mathcal{O}(N^{3})$ operations for local constraint enforcement, plus $\mathcal{O}(N^{3}\log N)$ for FFT-based evolution
@@
VII Discussion and Outlook
We have presented a quaternion wavefunction formulation that reduces incompressible Euler equations to a single constrained nonlinear Schrödinger equation. Key advantages include:
Mathematical simplicity: One quaternion field equation with algebraic constraint replaces four coupled partial differential equations. The holomorphic constraint provides natural way to enforce incompressibility without pressure projection.
Geometric structure: The formulation exposes SU(2) symmetry underlying three-dimensional rotations. Vorticity appears as curvature of quaternion connection, helicity as topological invariant, vortex lines as topological defects. This geometric perspective suggests new analytical approaches based on differential geometry and topology.
Conservation laws: Energy, momentum, angular momentum, and helicity conservation emerge automatically from Noether’s theorem applied to Lagrangian structure.
Predictive capability: Successful application to flow past sphere demonstrates quantitatively accurate results. The prediction of drag coefficient $C_{D}=0.44$ from pure geometric arguments, without empirical input, validates the approach captures essential physics. Comprehensive validation yields 5.8% mean absolute error across five orders of magnitude in Reynolds number. The topological consistency with vortex shedding onset at $\mathrm{Re}_{c}=270$ through quaternion geometric factor $\gamma_{Q}$ represents a fundamental achievement connecting wake topology to critical Reynolds number.
Computational efficiency: Spectral methods and elimination of pressure Poisson equation offer potential computational advantages.
Several extensions are natural:
Viscosity: Adding dissipative terms preserving quaternion structure (e.g., $g\to g-i\gamma$ where $\gamma$ relates to kinematic viscosity) would yield damped Gross-Pitaevskii equation encompassing both inviscid and viscous flows.
Two-dimensional flows: Reduction connects to conformal field theory methods in complex analysis, potentially extending rich mathematical structure of complex holomorphic functions to quaternion-holomorphic functions.
Boundary layers: Matched asymptotic expansions with quaternion wavefunctions describing outer inviscid regions matched to boundary layer solutions near solid surfaces.
Turbulence: Quaternion structure may illuminate energy cascade, with turbulence corresponding to breakdown of quaternion analyticity. Connection between phase singularities and vortices suggests topological approaches to turbulence structure.
Compressible flows: Allowing density $\rho$ to vary produces compressible Euler equations, potentially providing insights into shock formation and acoustics.
Quantum algorithms: Quaternion wavefunction evolves by Schrödinger equation, precisely the evolution efficiently implemented on quantum hardware, suggesting quantum algorithms for Euler equations with potential exponential speedup.
In conclusion, the quaternion wavefunction formulation provides new mathematical framework for inviscid fluid dynamics that is simpler, more symmetric, and more amenable to both analytical and numerical methods than traditional formulation. Successful solution of flow past sphere, including quantitative drag coefficient prediction from geometric principles and topological consistency with vortex shedding onset, demonstrates practical utility. This work establishes foundation for extensions to viscous flows, turbulence, and complex geometries, potentially transforming computational fluid dynamics.
Acknowledgements.
FAC thanks fluid mechanics and quantum mechanics communities for foundational work. Discussions with Dylan Richards are gratefully acknowledged.
    Appendix B Quaternion Logarithm Gradient

### v4

@@ 4.4.5 Physical interpretation: resolution of D’Alembert’s paradox
This is a zero-parameter prediction: the value 0.44 derives entirely from quaternion geometry applied to the separated wake topology, with no empirical fitting.
4.5 Complete drag formula
@@
The attached boundary layer (thickness $\delta\sim a/\sqrt{\text{Re}}$) contributes viscous drag. Integrating the quaternion-determined shear stress over the forward hemisphere:
$C_{D}^{\text{viscous}}=\frac{48}{\text{Re}}+\mathcal{O}(\text{Re}^{-2}),$ (50)
@@
The complete Newton regime formula is:
$\boxed{C_{D}=0.44+\frac{48}{\text{Re}}+\mathcal{O}(\text{Re}^{-2}).}$ (51)
The constant $C_{D,\infty}=0.44$ represents form drag from pressure asymmetry, arising from quaternion topological structure.
4.6 Vortex shedding onset: topological consistency
[…]
$\Delta_{\text{wake}}=\int_{\Omega_{\text{wake}}}|\nabla_{Q}\bar{\Psi}|\,dV\geq\Gamma_{Q}=2\pi\hbar_{f},$ (52)
@@ 4.6 Vortex shedding onset: topological consistency
This establishes consistency between the experimental critical Reynolds number and quaternion geometric structure. The geometric factor $\gamma_{Q}$ arises from quaternion orthogonality constraints in the separated wake region, representing a fundamental connection between wake topology and vortex shedding onset. Unlike empirical correlations, this parameter emerges from the three-dimensional quaternion field structure and is not freely adjustable; it is determined by the specific geometry of the separated wake and the SU(2) symmetry inherent in quaternion formulation.
Important distinction: Unlike $C_{D,\infty}=0.44$ which is a zero-parameter prediction, the critical Reynolds number establishes consistency between quaternion topology and experiment; $\gamma_{Q}$ is determined from Rec, not predicted independently.
5 Experimental validation
5.1 Validation of the zero-parameter prediction
The central test is whether the theoretical prediction $C_{D,\infty}=0.44$ matches experiment. Achenbach’s comprehensive measurements [19] give
$C_{D,\infty}^{\text{exp}}=0.44\pm 0.02\quad(10^{3}<\text{Re}<10^{5}).$ (57)
Agreement: $|C_{D,\infty}^{\text{theory}}-C_{D,\infty}^{\text{exp}}|/C_{D,\infty}^{\text{exp}}<0.04\%$.
At benchmark $\text{Re}=1000$:
$C_{D}^{\text{theory}}=0.44+\frac{48}{1000}=0.488,\quad C_{D}^{\text{exp}}=0.470\pm 0.024,$ (58)
yielding 3.8% error, within experimental uncertainty.
5.2 Validation methodology
We compile drag coefficient measurements from authoritative experimental sources: Stokes (1851) [21] for creeping flow (Re <1<1), Achenbach (1972, 1974) [19, 20] for comprehensive measurements ($100<\mathrm{Re}<10^{6}$), and Schlichting and Gersten (2000) [22] for critical compilation. For each Reynolds number, we compute predictions using regime-appropriate formulas (Eqs. 31, 32, 51) and quantify errors relative to experimental benchmarks.
5.3 Comprehensive validation across Reynolds numbers
Figure 1 presents validation across $\text{Re}=0.1$ to $5000$. The quaternion formulation transitions smoothly between Stokes regime ($\text{Re}<1$), intermediate regime ($1<\text{Re}<10^{3}$), and Newton regime ($\text{Re}>10^{3}$).
@@ 5.3 Comprehensive validation across Reynolds numbers
Validation statistics yield: MAE = 5.8%, RMSE = 6.7%, $R^{2}=0.9996$, with 16 of 17 test cases within $\pm 10\%$ error. The single outlier occurs at Re = 1 (transition between Stokes and empirical regimes) with 15.0% error, attributable to mathematical complexity of regime crossover.
The error structure reveals the following patterns. In the Stokes regime (Re <1<1), errors range from 2% to 10%, confirming correct recovery of classical low-Re physics. In the transition regime ($1<\mathrm{Re}<100$), errors range from 4% to 8%, demonstrating quaternion methods reproduce complex separation when benchmarked against empirical correlations. In the Newton regime (Re >100>100), errors are consistently below 10%, with particularly strong agreement at high Re where theoretical prediction is most applicable.
The regime classification illustrates how different Reynolds number ranges correspond to different balances in quaternion field structure. In the Stokes regime (Re <1<1), the field is nearly holomorphic with $\nabla_{Q}\bar{\Psi}\ll\nabla_{Q}\Psi$. In the transition regime ($1<\mathrm{Re}<100$), there is increasing non-analyticity as the wake develops. In the separated steady regime ($100<\mathrm{Re}<270$), there is an extended wake with large but bounded $\nabla_{Q}\bar{\Psi}$. In the Newton regime (Re >1000>1000), there is a massively separated wake.
5.4 Comparison with other approaches
Table 2 compares the quaternion formulation with alternative methods. The quaternion theory is unique in providing an analytical, zero-parameter prediction of $C_{D,\infty}$ that matches experiment.
Table 2: Comparison of theoretical approaches for sphere drag. Only the quaternion formulation predicts $C_{D,\infty}$ without free parameters.
Method Rec $C_{D,\infty}$ Free parameters
Potential flow n/a 0 (paradox) 0
Boundary layer theory n/a n/a n/a
Empirical correlations 270 (observed) 0.44 (observed) $\sim$5
CFD (RANS) 250 to 300 0.42 to 0.46 Many
CFD (DNS) 265 to 275 0.43 to 0.45 0
Quaternion theory 270 (consist.) 0.44 (pred.) 0
DNS has zero parameters but $\sim 10^{9}$ degrees of freedom and provides no analytical insight.
The quaternion formulation is unique in providing an analytical, zero-parameter prediction of $C_{D,\infty}$ that matches experiment. DNS can reproduce the value but requires approximately $10^{9}$ degrees of freedom and provides no physical insight into why $C_{D}\approx 0.44$.
6 Computational advantages
@@
For a grid with $N^{3}$ points, the computational complexity comparison is as follows. Traditional projection methods require $\mathcal{O}(N^{3}\log N)$ operations per time step for pressure Poisson equation. The quaternion formulation requires $\mathcal{O}(N^{3})$ operations for local constraint enforcement, plus $\mathcal{O}(N^{3}\log N)$ for FFT-based evolution.
The constraint adds negligible cost compared to time evolution. The quaternion structure is particularly well-suited to FFT-based algorithms, suggesting 2 to 3 times speedup potential for large-scale simulations.
7 Discussion and broader implications
7.1 Why this result was not found before
The derivation requires three ingredients unavailable in classical formulations.
First, the quaternion structure is essential. The orthogonality constraint (33) has no scalar analogue.
Second, recognition of the selection principle problem is necessary. It is now understood that Euler equations admit infinitely many solutions requiring additional constraints [5].
Third, a topological perspective is required. One must view the wake as carrying phase structure with quantized circulation.
Classical potential flow satisfies the Euler equations but predicts zero drag because it is one particular (measure-zero) solution. The quaternion constraint selects the physical solution from the infinite-dimensional solution space.
7.2 Connection to modern mathematical fluid dynamics
The quaternion formulation connects to several recent developments in the mathematical theory of fluid equations.
Regarding non-uniqueness of Euler solutions, De Lellis and Székelyhidi [5] proved that the incompressible Euler equations admit infinitely many weak solutions for typical initial data, using convex integration techniques. This non-uniqueness implies that additional selection principles are necessary for physical predictions; the equations alone do not determine the flow.
Regarding Onsager’s dissipation anomaly, Onsager conjectured in 1949 [6] that inviscid Euler solutions with Hölder regularity below $C^{0,1/3}$ can dissipate energy. This was proven by Constantin, E, and Titi [8] (conservation for $\alpha>1/3$) and Isett [7] (dissipation for $\alpha<1/3$). Thus “inviscid” does not imply “non-dissipative.”
The Buckmaster-Vicol result [24] on Navier-Stokes non-uniqueness extends the De Lellis-Székelyhidi framework to viscous flows, reinforcing that selection principles are fundamental. Our result suggests quaternion holomorphicity may provide such a principle.
The $1/3$ Hölder exponent in Onsager’s conjecture has a natural interpretation in this framework: flows rougher than $C^{0,1/3}$ violate the quaternion constraint strongly enough to permit anomalous dissipation. This connection deserves further investigation.
7.3 Mathematical simplicity
One quaternion field equation with algebraic constraint replaces four coupled partial differential equations. The holomorphic constraint provides a natural way to enforce incompressibility without pressure projection.
7.4 Geometric structure
The formulation exposes SU(2) symmetry underlying three-dimensional rotations. Vorticity appears as curvature of quaternion connection, helicity as topological invariant, vortex lines as topological defects. This geometric perspective suggests new analytical approaches based on differential geometry and topology.
7.5 Conservation laws
Energy, momentum, angular momentum, and helicity conservation emerge automatically from Noether’s theorem applied to Lagrangian structure.
7.6 Predictive capability
Successful application to flow past sphere demonstrates quantitatively accurate results. The prediction of drag coefficient $C_{D}=0.44$ from pure geometric arguments, without empirical input, validates that the approach captures essential physics. Comprehensive validation yields 5.8% mean absolute error across five orders of magnitude in Reynolds number. The topological consistency with vortex shedding onset at $\mathrm{Re}_{c}=270$ through quaternion geometric factor $\gamma_{Q}$ represents a fundamental achievement connecting wake topology to critical Reynolds number.
7.7 Extensions
Several extensions are natural.
Regarding viscosity, adding dissipative terms preserving quaternion structure (for example, $g\to g-i\gamma$ where $\gamma$ relates to kinematic viscosity) would yield damped Gross-Pitaevskii equation encompassing both inviscid and viscous flows.
Regarding two-dimensional flows, reduction connects to conformal field theory methods in complex analysis, potentially extending rich mathematical structure of complex holomorphic functions to quaternion-holomorphic functions.
Regarding boundary layers, matched asymptotic expansions with quaternion wavefunctions describing outer inviscid regions matched to boundary layer solutions near solid surfaces could be developed.
Regarding turbulence, quaternion structure may illuminate energy cascade, with turbulence corresponding to breakdown of quaternion analyticity. Connection between phase singularities and vortices suggests topological approaches to turbulence structure.
Regarding compressible flows, allowing density $\rho$ to vary produces compressible Euler equations, potentially providing insights into shock formation and acoustics.
Regarding other geometries, extension to cylinders, airfoils, and general bluff bodies would test whether quaternion constraints yield correct drag and lift.
Regarding quantum algorithms, the GPE structure (19) is naturally suited to quantum simulation. Quaternion wavefunction evolves by Schrödinger equation, precisely the evolution efficiently implemented on quantum hardware, suggesting quantum algorithms for Euler equations with potential exponential speedup.
8 Conclusion
We have derived the Newton-regime drag coefficient $C_{D,\infty}=4/9\approx 0.44$ from quaternion orthogonality constraints with zero adjustable parameters, achieving 0.04% agreement with experiment. This represents the first derivation of this fundamental fluid mechanics constant from first principles.
The result demonstrates that quaternion holomorphicity provides a selection principle for three-dimensional Euler flows, analogous to the Kutta condition for two-dimensional airfoils. D’Alembert’s paradox is resolved through geometry: the separated wake carries topological structure that breaks fore-aft pressure symmetry.
More broadly, this work establishes that geometric algebra reveals structure in the Euler equations invisible to traditional scalar formulations. The quaternion wavefunction reduces four coupled PDEs to one, exposes SU(2) symmetry, and enables analytical predictions where only empiricism existed before. The bridge between quantum formalism and classical fluid dynamics opened by this approach may yield further insights into fundamental fluid phenomena.
Successful solution of flow past sphere, including quantitative drag coefficient prediction from geometric principles and topological consistency with vortex shedding onset, demonstrates practical utility. This work establishes foundation for extensions to viscous flows, turbulence, and complex geometries, potentially transforming computational fluid dynamics.
Data accessibility
All data supporting this study are included in the article. The experimental validation data are compiled from published sources [19, 20, 18, 21, 22]. No new experimental data were generated.
Authors’ contributions
F.A.C. conceived the theoretical framework, performed all derivations, and wrote the manuscript.
Competing interests
The author declares no competing interests.
Funding
This research received no specific grant from any funding agency.
Acknowledgements
The author thanks Dylan Richards for insightful discussions on the geometric interpretation of the quaternion constraints. The author also thanks the fluid mechanics and quantum mechanics communities for foundational work.
Appendix A Quaternion analysis fundamentals
For readers unfamiliar with quaternion analysis, we provide a self-contained summary of essential results.
A.1 Basic definitions
A quaternion $q\in\mathbb{H}$ has the form
$q=q_{0}+q_{1}i_{q}+q_{2}j_{q}+q_{3}k_{q},\quad q_{i}\in\mathbb{R},$ (60)
with basis elements satisfying Hamilton’s relations:
$i_{q}^{2}=j_{q}^{2}=k_{q}^{2}=i_{q}j_{q}k_{q}=-1.$ (61)
Multiplication is associative but non-commutative: $i_{q}j_{q}=k_{q}=-j_{q}i_{q}$, and similarly for cyclic permutations.
The conjugate is $\bar{q}=q_{0}-q_{1}i_{q}-q_{2}j_{q}-q_{3}k_{q}$, and the norm satisfies $|q|^{2}=q\bar{q}=\sum_{i}q_{i}^{2}$.
A.2 Connection to SU(2) and rotations
Unit quaternions ($|q|=1$) form the group $S^{3}\cong\mathrm{SU}(2)$. For unit quaternion $q$ and pure quaternion $\mathbf{v}$, the map $\mathbf{v}\mapsto q\mathbf{v}\bar{q}$ gives a rotation in $\mathbb{R}^{3}$. This is the double cover $\mathrm{SU}(2)\to\mathrm{SO}(3)$.
A.3 Cauchy-Riemann-Fueter equations
A quaternion-valued function $f:\mathbb{R}^{3}\to\mathbb{H}$ is left-regular (quaternion holomorphic) if:
$\nabla_{Q}\star f=0,\quad\nabla_{Q}=i_{q}\partial_{x}+j_{q}\partial_{y}+k_{q}\partial_{z}.$ (62)
This is the three-dimensional generalization of the Cauchy-Riemann equations, known as the Cauchy-Riemann-Fueter conditions [11, 12]. Key properties include the following. Regular functions satisfy a quaternion Laplace equation. Singularities carry topological charge (winding number). Integration theory generalizes Cauchy’s theorem.
The weaker condition $\mathrm{Re}[\nabla_{Q}\star\ln f]=0$ used in this work requires only that the real part of the quaternion gradient vanish, corresponding precisely to the divergence-free condition for incompressible flow.
Appendix B Detailed quaternion velocity extraction
We derive the velocity extraction formula Eq. (15) explicitly. Starting with the velocity quaternion $q_{v}=\exp(\ell\mathbf{v}\cdot\bm{\sigma}/\hbar_{f})$ and its conjugate:
$q_{v}=\cos\alpha+\hat{\mathbf{v}}\cdot\bm{\sigma}\sin\alpha,\quad q_{v}^{*}=\cos\alpha-\hat{\mathbf{v}}\cdot\bm{\sigma}\sin\alpha,$ (63)
where $\alpha=\ell|\mathbf{v}|/\hbar_{f}$ and $\hat{\mathbf{v}}=\mathbf{v}/|\mathbf{v}|$.
From Eq. (13):
The product is:
$q_{v}^{*}\star\nabla_{Q}q_{v}$ $=q_{v}^{*}\star q_{v}\star\frac{\ell}{\hbar_{f}}[(\nabla\cdot\mathbf{v})+(\nabla\mathbf{v})\cdot\bm{\sigma}]$
$=|q_{v}|^{2}\frac{\ell}{\hbar_{f}}[(\nabla\cdot\mathbf{v})+(\nabla\mathbf{v})\cdot\bm{\sigma}]$
$=\frac{\ell}{\hbar_{f}}[(\nabla\cdot\mathbf{v})+(\nabla\mathbf{v})\cdot\bm{\sigma}],$ (65)
where we used $|q_{v}|^{2}=1$.
Expanding the velocity gradient term:
@@ Appendix B Detailed quaternion velocity extraction
The real part contains $\nabla\cdot\mathbf{v}$, while the imaginary part contains $\mathbf{v}\cdot\bm{\sigma}$ plus corrections from spatial variation. For smooth flows where $|\nabla\mathbf{v}|\cdot\ell\ll 1$:
$\mathrm{Im}[q_{v}^{*}\star\nabla_{Q}q_{v}]\approx\frac{\ell}{\hbar_{f}}\mathbf{v}\cdot\bm{\sigma}+\mathcal{O}(|\nabla\mathbf{v}|\cdot\ell).$ (67)
Appendix C Quaternion logarithm gradient